given tan0=2 and sec0=square root of 5, determine sin0
Answer by lwsshak3(11628)

given tan0=2 and sec0=square root of 5, determine sin0
use x for 0:
tanx=2
secx=√5
cosx=1/secx=1/√5=√5/5
tanx=sinx/cos2x=2
sinx=2cosx=(2√5)/5